# Towards a Unified Language Model for Knowledge-Intensive Tasks Utilizing External Corpus

### v1

@@ 4.3.1. Ablations for Retrieval Tasks
(1) Removing DocID understanding tasks reduces retrieval performance across all datasets for both T5 and LLama2-based CorpusLM. Specifically, we see notable decreases on the NQ, HoPo, and ELI5 datasets by $5.37\%$, $5.53\%$, and $18.42\%$ for CorpusLM (T5), respectively. For CorpusLM (Llama2), we also see notable declines on the WnWi and ELI5 datasets, with drops of $4.70\%$ and $7.40\%$, respectively. This underlines the crucial role of DocID understanding in enhancing CorpusLM’s retrieval capability. (2) Changing the ranking-oriented decoding to the traditional beam decoding leads to lower R-Precision in most datasets. Notably, for CorpusLM (T5), TQA and ELI5 saw reductions of $3.84\%$ and $7.74\%$ in R-Precision, highlighting the importance of this strategy for retrieval. For CorpusLM (Llama2), we also observe a consistent performance drop without the DocID understanding tasks, Moreover, removing the ranking-oriented decoding caused a more significant performance drop than that in CorpusLM (T5), particularly for the WnWi, ELI5, and WoW datasets, where we observed a decline of $7.05\%$, $16.34\%$, and $6.88\%$. This highlights the importance of our proposed ranking-oriented decoding strategy for decoder-only large language models like Llama2.
@@ 4.3.2. Ablations for Downstream Tasks
(1) Removing reference decoding, there’s a performance decline across all tasks. This is particularly pronounced on NQ and HoPo datasets, with a decrease of $2.92\%$ and $4.54\%$, respectively. This indicates that reference decoding plays a critical role in filtering out irrelevant information and generating more accurate answers. For CorpusLM (Llama2), removing reference decoding notably affects the accuracy in the Open Domain QA domain, especially for TQA and ELI5, with reductions of $2.85\%$ and $1.85\%$, respectively. This further enforces the value of a focused decoding strategy for high-quality answer generation. (2) Without noise sampling, we also observe a reduction in most datasets, with decreases of $0.61\%$ and $1.44\%$ on AY2 and on zsRE for T5-based CorpusLM, respectively. This suggests that introducing noise during training helps improve RAG performance. (3) The removal of continuous decoding leads to drops in most datasets for CorpusLM (T5), with a decline of $2.14\%$ on WnWi and $1.12\%$ on TQA, showcasing the importance of a streamlined decoding process.
@@ 4.4. Analysis of Ranking Capability
For the four datasets in the figure, we observe that Dense Retriever is consistently outperformed by Generative Retriever for Recall@1, but it surpasses GR at Recall@10. Notably, on the FEVER and NQ datasets, DR lags behind GR by 7.41% and 7.81% at Recall@1, yet it leads by 3.01% and 11.44% at Recall@10, respectively. This trend suggests that traditional GR methods are firmly focused on the top-1 ranked DocID, often at the expense of accuracy for other rop-ranked DocIDs, which is consistent with (Zhou et al., 2023a). Our CorpusLM, however, demonstrates superior performance to GR at all levels of Recall. With improvements of 3.68% and -10.94% in Recall@10 for FEVER and NQ, our model showcases an advanced capability for ranking enhancement. Moreover, CorpusLM consistently outpaces the DR at Recall@5 and is competitive at Recall@10, evidenced by margins of 1.98% higher and 0.45% lower than DR on the T-REx and NQ datasets, respectively. These results underline the strength of the ranking strategy employed by CorpusLM, leading to improved ranking performance.
@@
Method Parameters Memory Latency
RAG 626M
59.3G
106.7ms
MT-DPR+FID 440M
51.2G
160.9ms
[…]
  w/ Pipeline Decode 220M
426.1M
102.9ms
@@ 4.5. Efficiency Analysis for RAG Tasks
We then analyze the efficiency of our continuous RAG strategy, evaluating model parameters, memory footprint and query latency, as shown in Table 6. Our CorpusLM (T5) showcases a significant reduction in the number of parameters, memory usage, and query latency compared to other methods such as RAG and MT-DPR+FID. Specifically, the CorpusLM model has approximately 2.8-fold fewer parameters than the RAG (Lewis et al., 2020b) model, resulting in a substantially lower computational cost. Notably, it implies a 139.1-fold decrease in memory footprint, primarily due to the model only requiring storage for titles and sections as DocIDs rather than a large-scale dense document index. Moreover, compared with MT-DPR+FID (Maillard et al., 2021a), the CorpusLM model exhibits a 2.1-fold reduction in query latency. The table also highlights that even when utilizing a pipeline decode approach, our model maintains its efficiency advantage, with a 1.3-fold faster decoding speed compared to the multi-round input-output RAG method. Our CorpusLM model utilizes continuous decoding to combine generative retrieval and answer generation into a single step, which can accelerate response times and cut down on the computational costs of repeated interaction cycles in traditional RAG approaches.
5. Related Work
To enhance the performance of RAG models, various strategies have been developed, which involve training the retriever and reader modules separately using dense retrieval methods, such as DPR (Karpukhin et al., 2020), MT-DPR (Maillard et al., 2021b), DRQA (Chen et al., 2017) and FID (Izacard and Grave, 2020). Other strategies focus on joint training of both modules, either by updating the query encoder while keeping the document index static (Lewis et al., 2020b), or by updating both the query and document encoders and asymptotically updating the document index during training (Singh et al., 2021; Sachan et al., 2021; Izacard et al., 2023). Based on generative retrieval methods, SEAL (Bevilacqua et al., 2022), CorpusBrain (Chen et al., 2022) and UGR (Chen et al., 2023) train the GR models and also utilize separate generators to accomplish downstream KILT tasks. Recently, UniGen (Li et al., 2023c) introduced a unified framework for concurrent learning of GR and QA tasks. However, it uses separate decoders to generate DocIDs and answers, restricting its ability to generalize and scale.
@@ 6. Conclusion and Future Works
In this work, we propose CorpusLM, a unified language model that handles various knowledge-intensive tasks by integrating generative retrieval, closed-book generation, and RAG. By unifying all tasks and incorporating auxiliary DocID understanding tasks, we facilitate joint learning of retrieval and generation, thus enhance the model’s performance in KI tasks. Furthermore, our model is enhanced with two well-designed generation strategies: (1) a ranking-oriented DocID list generation strategy, which refines GR by directly learning from a DocID ranking list, to improve retrieval quality. (2) a continuous DocIDs-References-Answer generation strategy, which facilitates effective and efficient RAG. These generation strategies enable the model to perform retrieval and RAG in a continuous auto-regressive decoding process. Experiments conducted on 11 datasets in the KILT benchmark have shown our CorpusLM’s superior performance in both retrieval and downstream knowledge-intensive tasks.
This work opens up the potential for unifying various IR tasks (especially the retrieval task that cannot be easily performed by LLMs) into a single generative language model. For future works, there are three potential ways to further enhance CorpusLM: (1) Integrate more IR tasks, such as recommendation, summarization, etc.; (2) Develop a multi-modal, multi-lingual CorpusLM; (3) Following (Schick et al., 2023; Shen et al., 2023), it’s possible to employ generative retrieval as an internal ”tool”. This would enable CorpusLM to determine the appropriate instances to generate DocIDs, which could increase the factual accuracy of the generated text, without relying on external retrieval modules.

### v2

@@ 5.3.1. Ablations for Retrieval Tasks
(1) Removing DocID understanding tasks reduces retrieval performance across all datasets for both T5 and LLama2-based CorpusLM. Specifically, we see notable decreases on the NQ, HoPo, and ELI5 datasets by $5.37\%$, $5.53\%$, and $18.42\%$ for CorpusLM (T5), respectively. For CorpusLM (Llama2), we also see notable declines on the WnWi and ELI5 datasets, with drops of $4.70\%$ and $7.40\%$, respectively. This underlines the crucial role of DocID understanding in enhancing CorpusLM’s retrieval capability. (2) Changing the ranking-oriented decoding to the traditional beam decoding leads to lower R-Precision in most datasets. Notably, for CorpusLM (T5), TQA and ELI5 saw reductions of $3.84\%$ and $7.74\%$ in R-Precision, highlighting the importance of this strategy for retrieval. For CorpusLM (Llama2), we also observe a consistent performance drop without the DocID understanding tasks,
@@ 5.3.2. Ablations for Downstream Tasks
(1) Removing reference decoding, there’s a performance decline across all tasks. This is particularly pronounced on NQ and HoPo datasets, with a decrease of $2.92\%$ and $4.54\%$, respectively. For CorpusLM (Llama2), removing reference decoding notably affects the accuracy in the Open Domain QA domain, especially for TQA and ELI5, with reductions of $2.85\%$ and $1.85\%$, respectively. This indicates that reference decoding plays a critical role in filtering out irrelevant information and generating more accurate answers. (2) Without noise sampling, we also observe a reduction in most datasets, with decreases of $0.61\%$ and $1.44\%$ on AY2 and on zsRE for T5-based CorpusLM, respectively. This suggests that introducing noise during training helps improve RAG performance. (3) The removal of continuous decoding leads to drops in most datasets for CorpusLM (T5), with a decline of $2.14\%$ on WnWi and $1.12\%$ on TQA, showcasing the importance of a streamlined decoding process.
@@ 5.4. Analysis of Ranking Capability
For the four datasets in the figure, we observe that Dense Retriever is consistently outperformed by Generative Retriever for Recall@1, but it surpasses GR at Recall@10. Notably, on the FEVER and NQ datasets, DR lags behind GR by 7.41% and 7.81% at Recall@1, yet it leads by 3.01% and 11.44% at Recall@10, respectively. This trend suggests that traditional GR methods are firmly focused on the top-1 ranked DocID, often at the expense of accuracy for other rop-ranked DocIDs. Our CorpusLM, however, demonstrates superior performance to GR at all levels of Recall. With improvements of 3.68% and -10.94% in Recall@10 for FEVER and NQ, our model showcases an advanced capability for ranking enhancement. Moreover, CorpusLM consistently outpaces the DR at Recall@5 and is competitive at Recall@10. These results underline the strength of the ranking strategy employed by CorpusLM, leading to improved ranking performance.
@@
Method Parameters Storage Latency
RAG 626M 59.3G 106.7ms
MT-DPR+FID 440M 51.2G 160.9ms
CorpusLM (T5) 220M 426.1M 78.4ms
w/ Pipeline Decode 220M 426.1M 102.9ms
@@ 5.5. Efficiency Analysis for RAG Tasks
We then analyze the efficiency of our continuous RAG strategy, evaluating model parameters, memory footprint and query latency, as shown in Table 6. Our CorpusLM (T5) showcases a significant reduction in the number of parameters, memory usage, and query latency compared to other methods such as RAG and MT-DPR+FID. Specifically, the CorpusLM model has approximately 2.8-fold fewer parameters than the RAG (Lewis et al., 2020b) model, resulting in a substantially lower computational cost. Notably, it implies a 139.1-fold decrease in memory footprint, primarily due to the model only requiring storage for titles and sections as DocIDs rather than a large-scale dense document index. Moreover, compared with MT-DPR+FID (Maillard et al., 2021b), the CorpusLM model exhibits a 2.1-fold reduction in query latency. The table also highlights that even when utilizing a pipeline decode approach, our model maintains its efficiency advantage, with a 1.3-fold faster decoding speed compared to the multi-round input-output RAG method. Our CorpusLM model utilizes continuous decoding to combine generative retrieval and answer generation into a single step, which can accelerate response times and cut down on the computational costs of repeated interaction cycles in traditional RAG approaches.
@@ 6. Conclusion and Future Works
In this work, we propose CorpusLM, a unified language model that handles various knowledge-intensive tasks by integrating generative retrieval, closed-book generation, and RAG. By unifying all tasks and incorporating auxiliary DocID understanding tasks, we facilitate joint learning of retrieval and generation, thus enhance the model’s performance in KI tasks. This work opens up the potential for unifying various IR tasks (especially the retrieval task that cannot be easily performed by LLMs) into a single generative language model. Furthermore, our model is enhanced with two well-designed generation strategies: (1) a ranking-oriented DocID list generation strategy, which refines GR by directly learning from a DocID ranking list, to improve retrieval quality. (2) a continuous DocIDs-References-Answer generation strategy, which facilitates effective and efficient RAG. These generation strategies enable the model to perform retrieval and RAG in a continuous auto-regressive decoding process. Experiments conducted on 11 datasets in the KILT benchmark have shown our CorpusLM’s superior performance in both retrieval and downstream tasks.
Acknowledgements.
Zhicheng Dou is the corresponding author. This work was supported by Beijing Natural Science Foundation No. L233008, National Natural Science Foundation of China No. 62272467, the fund for building world-class universities (disciplines) of Renmin University of China, and Public Computing Cloud, Renmin University of China. The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE, and Beijing Key Laboratory of Big Data Management and Analysis Methods.